# Backstepping for Partial Differential Equations

### v2

@@ 5.2 Applications of Moving Boundary Domain Designs
The moving boundary domain problems described in Section 4.6 arise in various applications, including 3-D printing, estimation of polar ice, and crystal growth processes. In these applications, the interface between different phases (e.g., solid and liquid, ice and water) is dynamic, and its position plays a crucial role in determining the system’s overall behavior. In [173], the authors address the problem of regulating temperature in 3-D printers based on selective laser sintering. The dynamics of the melt pool is described by a moving boundary problem, where the interface between the solid and liquid phases evolves according to the heat input from the laser. The work [172] proposes a laser sintering control design for metal additive manufacturing by PDE backstepping. The objective is to stabilize the depth of the melt pool at a desired position. The authors utilize a 1-D approximation of the Stefan problem in the vertical direction, incorporating the in-domain effect of laser power; these results are robust even when applied to a more complex two-phase Stefan model and in the presence of measurement bias. On the other hand, [169] deals with the challenging problem of estimating the thickness of polar ice sheets, which involves a moving boundary due to the melting and freezing processes. The authors employ a backstepping observer that utilizes measurements of the ice surface elevation to estimate the ice thickness, enabling the monitoring of changes in polar ice volume and its impact on sea level rise. Another example includes [71], which addresses nonlinear stabilization through wave PDE dynamics with a moving uncontrolled boundary, with applications to oil drilling where the domain length changes with time and depends on the drill bit speed. Finally, in [105], the authors present an alternative design and implementation of a backstepping controller for regulating temperature in 3-D printers based on selective laser sintering which aims to achieve rapid and accurate temperature regulation, ensuring proper layer deposition and achieving desired print quality. The authors demonstrate the effectiveness of their approach through experimental validation on a real 3-D printer.
@@ 5.3 Flow Control
Flow control, encompassing the manipulation of fluid flows to achieve desired behaviors, represents a fertile ground for applications of PDE backstepping. Backstepping has enabled the design of boundary controllers and observers for a variety of flow models, including the Navier-Stokes equations, magnetohydrodynamic (MHD) systems, multi-layer flows, and pipe flows, offering new avenues for enhancing flow stability, reducing drag, and mitigating undesired flow phenomena. One of the earliest successes of PDE backstepping in flow control was the stabilization of the linearized 2-D Navier-Stokes Poiseuille flow using a closed-form feedback controller [267]. This work demonstrated the feasibility of designing explicit boundary control laws for a simplified flow model, paving the way for more complex scenarios such as transitions between profiles [279]. Subsequent efforts, such as [80], focused on motion planning and trajectory tracking for three-dimensional Poiseuille flow (which result in the well-known Orr-Sommerfeld-Squire equations). By combining backstepping with a motion planning algorithm, the authors designed a controller that drives the velocity field of the flow to a desired trajectory, illustrating the potential for achieving precise flow manipulation. The control of MHD channel flows, involving electrically conducting fluids subject to magnetic fields, has also benefited from backstepping designs. In [295], the authors addressed the stabilization of the linearized 2-D MHD channel flow using backstepping boundary control. This work highlighted the applicability of backstepping to more complex flow models, incorporating the interaction between fluid dynamics and electromagnetism. Further explorations of MHD state estimation with boundary sensors were presented in [278]. A common theme in these flow control applications is the utilization of spatial invariance to simplify the control design (see Section 4.3; techniques like the Fourier transform can be employed to convert the PDE system into an ensemble of parameterized lower-dimensional systems. This allows for the application of backstepping to each wave number separately, simplifying the control design and analysis.
@@ 6.3 Machine Learning and PDE backstepping
In a recent study [303], backstepping controllers were compared to reinforcement learning controllers for addressing stop-and-go traffic congestion on a freeway segment. The study found that while reinforcement learning controllers can achieve similar performance to backstepping controllers with less computation time, they typically require one thousand episodes of iterative training on a simulation model. Conducting this training in a collision-free manner within real traffic is impractical for PDE-based models, although some successful implementations exist for simpler ODE-based traffic models [201]. However, even in these cases, convergence during training cannot be guaranteed. As a result, reinforcement learning controllers are not a straightforward or entirely secure replacement for model-based control in real traffic systems. However, it is possible to successfully combine backstepping with machine learning. A recent development [56] introduced a promising approach, involving a direct deep-neural network approximation of the backstepping controller for a transport equation with re-circulation. This model-based machine-learning backstepping controller offers established convergence guarantees using the DeepONet general approximation theorem, along with a Lyapunov analysis. Current progress in this area demonstrate the enormous potential of neural operators to effectively learn backstepping gains and controllers while maintaining stability guarantees. This is reflected in the latest works, which include designing neural operator-based observers and controllers for parabolic PDEs [182], extending the approach to more complex hyperbolic PDE systems [286], developing adaptive neural operator-based control designs [197], and even implementing gain scheduling techniques using neural operators [198].
@@ 7.2 In-domain actuation
Some early results concerning potential extensions to in-domain actuation are [262] and [287] in the parabolic case. These two results require very specific conditions on the shape of the actuation in the former, and specific domain lengths in the latter (similar to rational-dependence). An extension to observer design with measurements of weighted averages can also be found in [261]. For hyperbolic systems, early results can be found in [311] and [231]. In particular, the last result links with the Fredholm transformation results in Section 4.7. Seeing the recent advances in Fredholm transformation development, it is possible that more generic results in this are could be attainable soon.
@@ 7.3 Backstepping and stochastic systems
Mean-square exponential stabilization of a stochastic reaction-diffusion equation has been considered in [293]. The backstepping approach was used to design a nominal stabilizing controller that guarantees the mean-square exponential stability. This primer result suggests most results presented in this survey to design stabilizing controllers for interconnected ODE-PDE systems could be extended to design mean-square stabilizing controllers for the stochastic counterparts. However, stabilizing in expectation should not be the only objective, and one must make sure the variance remains bounded to ensure reliability and closed-loop robustness. In this context, the methodology developed in [282] for a hyperbolic PDE interconnected with a stochastic differential equation seems promising. Designing backstepping-based risk-averse methods that scale with many different risk measures is a challenging task for future research.
@@ 7.4 Towards parametrizable target systems
with an appropriate choice of and to guarantee closed-loop stability. It was shown in [232] that the target system (78) was always reachable from the original system (51). In the other hand, the Port-Hamiltonian approach (PHS) [200, 228] can provide a structured methodology to emphasize and leverage the intrinsic physical properties of the original system (such as passivity, dissipativity, and reversibility) therefore guiding appropriate selections of and in the target system (78). All in all, incorporating degrees of freedom into the design results in modular controllers that can ensure various trade-offs between performance specifications, such as robustness margins, sensitivity, and convergence rate. Although we can comprehend the qualitative impact of the tuning parameters, quantifying the performance of the resulting controllers in terms of performance specifications remains an area of ongoing research.
@@ 7.5 Multi-dimensional Systems
While Section 4.3 discusses existing results for higher-dimensional domains, most backstepping applications remain fundamentally one-dimensional. Current multi-dimensional results are primarily limited to special geometries—such as balls, disks, and boxes—where symmetry or separability allows reduction to one or more 1D systems. Even in these cases, the resulting kernel equations become significantly more complex, often requiring sophisticated tools from complex analysis and special functions theory. A general theory for multi-dimensional backstepping that does not rely on dimensional reduction or special geometric properties would represent a significant breakthrough. Such a theory would need to address several challenges: the formulation of appropriate target systems in multiple dimensions, the well-posedness of the resulting multi-dimensional kernel equations, and the development of practical methods for computing the resulting kernel functions. Success in this direction could dramatically expand the applicability of backstepping to important physical systems that are inherently multi-dimensional, such as fluid flows, electromagnetic fields, and complex thermal systems.

### v3

@@ 5.2 Applications of Moving Boundary Domain Designs
The moving boundary domain problems described in Section 4.6 arise in various applications, including 3-D printing, estimation of polar ice melting and freezing, and crystal growth processes. In these applications, the interface between different phases (e.g., solid and liquid, ice and water) is dynamic, and its position plays a crucial role in determining the system’s overall behavior. In [174], the authors address the problem of regulating temperature in 3-D printers based on selective laser sintering. The dynamics of the melt pool is described by a moving boundary problem, where the interface between the solid and liquid phases evolves according to the heat input from the laser. The work [173] proposes a laser sintering control design for metal additive manufacturing by PDE backstepping. The objective is to stabilize the depth of the melt pool at a desired position. The authors utilize a 1-D approximation of the Stefan problem in the vertical direction, incorporating the in-domain effect of laser power; these results are robust even when applied to a more complex two-phase Stefan model and in the presence of measurement bias. On the other hand, [170] deals with the challenging problem of estimating the thickness of polar ice sheets, which involves a moving boundary due to the melting and freezing processes. The authors employ a backstepping observer that utilizes measurements of the ice surface elevation to estimate the ice thickness, enabling the monitoring of changes in polar ice volume and its impact on sea level rise. Another example includes [72], which addresses nonlinear stabilization through wave PDE dynamics with a moving uncontrolled boundary, with applications to oil drilling where the domain length changes with time and depends on the drill bit speed. Finally, in [106], the authors present an alternative design and implementation of a backstepping controller for regulating temperature in 3-D printers based on selective laser sintering which aims to achieve rapid and accurate temperature regulation, ensuring proper layer deposition and achieving desired print quality. The authors demonstrate the effectiveness of their approach through experimental validation on a real 3-D printer.
@@ 5.3 Flow Control
Flow control, encompassing the manipulation of fluid flows to achieve desired behaviors, represents a fertile ground for applications of PDE backstepping. Backstepping has enabled the design of boundary controllers and observers for a variety of flow models, including the Navier-Stokes equations, magnetohydrodynamic (MHD) systems, multi-layer flows, and pipe flows, offering new avenues for enhancing flow stability, reducing drag, and mitigating undesired flow phenomena. One of the earliest successes of PDE backstepping in flow control was the stabilization of the linearized 2-D Navier-Stokes Poiseuille flow using a closed-form feedback controller [270]. This work demonstrated the feasibility of designing explicit boundary control laws for a simplified flow model, paving the way for more complex scenarios such as transitions between profiles [284]. Subsequent efforts, such as [81], focused on motion planning and trajectory tracking for three-dimensional Poiseuille flow (which result in the well-known Orr-Sommerfeld-Squire equations). By combining backstepping with a motion planning algorithm, the authors designed a controller that drives the velocity field of the flow to a desired trajectory, illustrating the potential for achieving precise flow manipulation. The control of MHD channel flows involving electrically conducting fluids subject to magnetic fields has also benefited from backstepping designs. In [299], the authors addressed the stabilization of the linearized 2-D MHD channel flow using backstepping boundary control. This work highlighted the applicability of backstepping to more complex flow models, incorporating the interaction between fluid dynamics and electromagnetism. Further explorations of MHD state estimation with boundary sensors were presented in [283]. A common theme in these flow control applications is the utilization of spatial invariance to simplify the control design (see Section 4.3; techniques like the Fourier transform can be employed to convert the PDE system into an ensemble of parameterized lower-dimensional systems. This allows for the application of backstepping to each wave number separately, simplifying the control design and analysis.
@@ 6.3 Machine Learning and PDE backstepping
In a recent study [307], backstepping controllers were compared to reinforcement learning controllers for addressing stop-and-go traffic congestion on a freeway segment. The study found that while reinforcement learning controllers can achieve similar performance to backstepping controllers with less computation time, they typically require one thousand episodes of iterative training on a simulation model. Conducting this training in a collision-free manner within real traffic is impractical for PDE-based models, although some successful implementations exist for simpler ODE-based traffic models [202]. However, even in these cases, convergence during training cannot be guaranteed. As a result, reinforcement learning controllers are not a straightforward or entirely secure replacement for model-based control in real traffic systems. However, it is possible to successfully combine backstepping with machine learning. A recent development [56] introduced a promising approach, involving a direct deep-neural network approximation of the backstepping controller for a transport equation with re-circulation. This model-based machine-learning backstepping controller offers established convergence guarantees using the DeepONet general approximation theorem, along with a Lyapunov analysis. Current progress in this area demonstrate the enormous potential of neural operators to effectively learn backstepping kernels (or alternatively, gains, as explained in [281]) while maintaining stability guarantees. This is reflected in the latest works, which include designing neural operator-based observers and controllers for parabolic PDEs [183, 238], extending the approach to more complex hyperbolic PDE systems [290], developing adaptive neural operator-based control designs [198], and even implementing gain scheduling techniques using neural operators [199].
@@ 7.2 In-domain actuation
Some early results concerning potential extensions to in-domain actuation are [264] and [291] in the parabolic case. These two results require very specific conditions on the shape of the actuation in the former and specific domain lengths in the latter (similar to rational dependence). An extension to observer design with measurements of weighted averages can also be found in [263]. For hyperbolic systems, early results can be found in [316] and [232]. In particular, the last result links with the Fredholm transformation results in Section 4.7. Seeing the recent advances in Fredholm transformation development, it is possible that more generic results in this area could be obtained soon.
@@ 7.3 Backstepping and stochastic systems
Mean-square exponential stabilization of a stochastic reaction-diffusion equation has been considered in [297]. The backstepping approach was used to design a nominal stabilizing controller that guarantees the mean-square exponential stability. This result suggests most designs presented in this survey to design stabilizing controllers for interconnected ODE-PDE systems could be extended to mean-square stabilizing controllers for the stochastic counterparts. However, stabilizing in expectation should not be the only objective, and one must make sure the variance remains bounded to ensure reliability and closed-loop robustness. In this context, the methodology developed in [286] for a hyperbolic PDE interconnected with a stochastic differential equation seems promising. Designing backstepping-based risk-averse methods that scale with many different risk measures is a challenging task for future research.
@@ 7.4 Towards parametrizable target systems
with an appropriate choice of and to guarantee closed-loop stability. It was shown in [233] that the target system (78) was always reachable from the original system (51). On the other hand, the Port-Hamiltonian approach (PHS) [201, 229] can provide a structured methodology to emphasize and leverage the intrinsic physical properties of the original system (such as passivity, dissipativity, and reversibility) therefore guiding appropriate selections of and in the target system (78). All in all, incorporating degrees of freedom into the design can result in modular controllers that allow for various trade-offs between performance specifications, such as robustness margins, sensitivity, and convergence rate. Although we may have an idea of the qualitative impact of tuning parameters, quantifying the performance of the resulting controllers in terms of performance specifications remains an area of ongoing research.
@@ 7.5 Multi-dimensional Systems
While Section 4.3 discusses existing results for higher-dimensional domains, most backstepping applications remain fundamentally one-dimensional. Current multi-dimensional results are primarily limited to special geometries—such as balls, disks, and boxes—where symmetry or separability allows reduction to one or more 1D systems. Even in these cases, the resulting kernel equations become significantly more complex, often requiring sophisticated tools from complex analysis and special functions theory. A general theory for multi-dimensional backstepping that does not rely on dimensional reduction or special geometric properties would represent a significant breakthrough. Such a theory would need to address several challenges: the formulation of appropriate target systems in multiple dimensions, the well-posedness of the resulting multi-dimensional kernel equations, and the development of practical methods for computing the resulting kernel functions. Success in this direction could dramatically expand the applicability of backstepping to important physical systems that are inherently multi-dimensional, such as fluid flows, electromagnetic fields, and complex thermal systems. Addressing some of these challenges, reference [266] explores these issues and investigates a potential direction using domain extension techniques, particularly for handling irregular domains.